# Extragalactic Stellar Tidal Streams: Observations meet Simulation

### v1

@@ 5 Stream Detection
Figure 11: Detection rate curves for the COCO, IllustrisTNG50 and Auriga cosmological simulations and the detection rate observed for the DES sample (black square indicates the dispersion in surface brightness limit of the DES sample and the dispersion in the detection rate according to a binomial distribution, see Miro-Carretero et al. (2024)). The percentage of the streams detected by visual inspection is plotted versus the surface brightness limit at which such detection was possible. The bars indicate the Poisson confidence interval.
@@ 5.1 Stellar Streams in the COCO Simulation
We have generated mock images with surface brightness limit between 25 and 34 mag arcsec-2 in intervals of 1 mag arcsec-2. As the depth of the image increases, that is, the fainter the surface brightness limit is, the clearer the streams appears to the observer, as can be seen in Figure 10.
@@
The result is a curve indicating the percentage of streams detected within the COCO halo sample as a function of the image depth, that is, as a function of the image surface brightness limit. The result is depicted in Figure 11. The curve follows an approximately linear trend between surface brightness limits of 26 and 34 mag arcsec-2, with a $12.5\%$ increase in streams detected per 1 mag arcsec-2 increase in surface brightness limit. In $97\%$ of the mock-images, we detect by visual inspection what we consider to be (part of) a stream up to a SB-limit of 34 mag arcsec-2. However, we estimate that that only 89% of such faint structures could actually be measured with a reasonable level of error in real images.
Regarding the stream morphology, in the COCO sample, at SB-limit 28-29 arcsec-2, around 70-80% of the detected streams are shells (a segment of the wider morphology class known as umbrella, see stream morphology classification in Miro-Carretero et al. (2024) while at SB-limit 34 arcsec-2 around 80-90% of the detected streams are shells, with only 2-3% of the streams displaying a circular morphology.
@@
Figure 13: Examples of IllustrisTNG50 mock-images TNG50-394621, TNG50-440407, TNG50-469487 and TNG50-472548 at different surface brightness limits. From left to right 29, 30 and 31 mag arcsec-2. The streams appear clearer as the depth of the image increases from left to right. All images are $10\times 10$ arcmin.
5.2 Stellar Streams in the IllustrisTNG50 Simulation
@@
We have transformed the surface brightness maps into images with counts and added sky background corresponding to 10 levels of surface brightness limit between 25 and 34 mag arcsec-2 ( see Section 4). As we increase the depth of the image (apply a higher background noise level to the image corresponding to a fainter surface brightness limit), streams become more visible and can be detected by visual inspection, as can be seen in Figure 13. We have visually inspected the resulting $\sim 600$ images and for each halo / host galaxy, the surface brightness limit at which a streams could first be visually appreciated was identified and noted. As with our analysis of COCO, FITS images were inspected with DS9.
Figure 14: Stellar streams detection rate as a function of the image surface brightness limit for the IllustrisTNG50 simulation. The percentage of the streams detected by visual inspection is plotted versus the surface brightness limit at which such detection was possible. The solid line corresponds to a stellar mass range of M/M = 10.48 and log M/M = 10.76, while the dashed line corresponds to an extended mass range up to log M/M = 10.9.
@@ 5.3 Stellar Streams in the Auriga Simulation
We have available 30 Auriga zoom simulations of Milky Way-mass halos at $z<0.02$. Since this sample may not be statistically representative, and in order to compare with the COCO and IllustrisTNG50 simulations, three surface brightness maps have been generated from each halo, taking the three orthogonal axes of the simulation coordinate system as lines of sight. None of the axes are aligned with face-on or edge-on directions so that overall the orientation of the galaxies is random. This process results in 90 images, each of which we combine with a range of sky backgrounds between the SB limit of 25 and 34 mag arcsec-2, at intervals of 1 mag arcsec-2, thus yielding 900 mock images in total. Examples of halos with streams of different morphology can be seen in Figure 15.
@@
Figure 17: Comparison of the stellar stream detection rate as a function of the image surface brightness limit for IllustrisTNG50 (green), COCO (red) and AURIGA (blue) with the results of observations. (1) The calculation of the SB limit follows a different method (2) The detection rate has been derived for all LSB features, including streams.
Figure 11 shows the detection rate curve obtained for Auriga from visual inspection of the mock images, together with those obtained for COCO and IllustrisTNG50, and the observational results of the DES sample. The figure shows the percentage of the sample for which at least one stream is detected for each of the 10 surface brightness limit levels analysed. The curve indicates detections of streams with a reasonable level of confidence. When we detect a LSB feature but have doubts about whether that LSB feature constitutes a stream, we take the conservative approach of not counting it towards the results included in the figureTwo of the halos (Au20 and Au30) seem to show an ongoing major merger between two galaxies of similar apparent size, clearly visible at a surface brightness limit of 27 mag arcsec-2. These two halos have been discarded, because we are looking for remnants of minor mergers only. LSB structures that could appear similar to streams at fainter surface brightness limits are not accounted for in the detectability curve.
@@
Table 2: Percentages of observed main stream morphologies shell and circle in the the DES sample and in the mock images for the COCO, IllustrisTNG50 and AURIGA simulations.The 4th column indicates the total number of streams detected in the sample. Note that some images present more than one stream.
@@ 6 Discussion
We now compare the predictions of the cosmological simulations with one another and with the observational data regarding frequency, detectability and morphology of streams. Looking at Figure 17, the cosmological simulations all seem to predict that, for a surface brightness limit of 32 $\mathrm{mag\,arcsec^{-2}}$, almost 70% of galaxies in the mass range we study have one or more detectable streams.
However, the simulations show discrepancies with one another regarding detection rates for surface brightness limits brighter than $32\,\mathrm{mag\,arcsec^{-2}}$. In particular, the two hydrodynamical simulations, IllustrisTNG50 and Auriga, agree well with each other but predict a lower detectability rate than COCO. For limits fainter than $32\,\mathrm{mag\,arcsec^{-2}}$, however, IllustrisTNG50 has a lower detection rate than COCO and Auriga (which agree well with each other, in this regime). The simulations also show discrepancies with one another regarding the morphology of the streams detected, as will be discussed further down in this section.
@@
As discussed in (Cooper et al., 2017), when comparing particle tagging simulations with hydrodynamical methods, even with identical initial conditions, it can be difficult to separate effects due to the dynamical approximation of particle tagging from the effects of different star formation models. In our case, the IllustrisTNG and COCO models have both been calibrated to observations of the galaxy mass and luminosity functions at $z=0$ (in the case of COCO via the semi-analytic model of Lacey et al., 2016), as well as other low-redshift data, notably the galaxy size–mass relation. The fundamental relationship between stellar mass and virial mass in both simulations agrees well with (for example) that inferred from galaxy abundance matching. Although the typical star formation histories of host galaxies in our sample, and their stream progenitors, may still differ in detail between the two simulations, we expect that they are broadly similar. This makes it more likely (although by no means certain) that the differences we observe are related to dynamical factors, rather than differences in how stars populate dark matter halos.
@@
Cooper et al. (2017) explore the consequences of neglecting these factors when predicting satellite disruption (and hence stream and stellar halo formation) in particle tagging models. In the hydrodynamical model used in that work, massive satellites were found to disrupt somewhat earlier than realizations of the same systems in a particle tagging model [^7] This of course depends on the detail of the hydrodynamical scheme and its subgrid recipes for star formation and feedback; the specific scheme used in IllustrisTNG was not examined by Cooper et al. (2017). A detailed case study of a single bright stream relevant to this discussion is given in Appendix A of that paper.. This could explain why more streams are visible at surface brightness limits of $\lesssim 28\,\mathrm{magarcsec^{-2}}$ in COCO compared to IllustrisTNG; if satellites are disrupted earlier in IllustrisTNG, their streams may have more time to phase mix, lowering their surface brightness. A similar argument could also explain the greater abundance of shells detected in COCO. Shells originate from satellites on radial orbits (Newberg & Carlin, 2016). The relative fraction of radial and circular orbits may differ between IllustrisTNG and COCO, because the evolution of the progenitor’s orbit during pericentric passages (due to exchange of angular momentum with the host) may be significantly different with and without a massive baryonic disk. A related effect is explored by Valenzuela & Remus (2024), who study the relation between the rotation of the hosts and the presence and morphology of tidal features One of the conclusions of that work is that shells appear more frequently in slow rotating hosts than in fast rotating hosts. This hypothesis could be tested in future work. More generally, further exploration of these differences particle tagging and hydrodynamical simulations could help to understand how observations of streams can constrain the nature of the gravitational potentials of stream progenitors and their host galaxies.
@@
Turning now to comparison between the simulations and the observational data, $8.7\pm 1.1\%$ of galaxies in the DES sample have detectable streams at an average $r$-band surface brightness limit of 28.65 mag arcsec-2. of 28.65 mag arcsec-2 is 8.7 $\pm$ 1.1 %. This seems to match very well with the predictions of IllustrisTNG50 and Auriga, and is lower than the prediction from COCO.
@@
Stream morphology is not an observable that can be used reliably to constrain simulations, because it is strongly dependent on the line of sight along which the stream is observed. Nevertheless, comparison of the fraction of streams with different morphologies is interesting because it could be related to the same dynamical differences between simulations methods that give rise to differences in stream abundance. 20-43% of streams in the IllustrisTNG50 sample are shells, which is within the range found in the DES sample, and 10-16% have circular morphology, which is also not too far from the observations. As noted above, there is a significant discrepancy in the relative fractions of different stream morphologies between observations and the COCO simulations. In the COCO mock images, 70-90% of streams are identified as having Umbrella/Shell morphology, while in the DES sample, only 27-38% of streams are identified as such. The COCO mock images show only 2-3 % of streams with circular morphologies, compared to 21-35 % in the DES images. Auriga predicts 41-62% of streams with Umbrella/Shell morphologies, significantly above the fraction in the DES sample, and 10-24% of stream with circular morphology, somewhat below the DES figures. Table 2 summarises the stream morphology findings for the different samples derived from the total number of streams detected. Note that more than one stream is detected in some of the halos. It is beyond the scope of this work to analyse the reasons for these discrepancies in stream morphology; as noted above, further work would be a valuable contribution to understanding the relationship between the observable properties of streams and the galaxy formation process as a whole.
@@ 6.1 Comparison with Previous Work
  • Miskolczi et al. (2011) reports a frequency of 6-19% at a SB-limit of $\sim$ 27 mag $arcsec^{-}2$. We consider the lowest value of 6% as the most reliable for maximising purity in the detection of streams (see 17.
@@
  • Morales et al. (2018) reports a frequency of $\sim$ 10% at a SB limit of 28 mag $arcsec^{-}2$.
  • Bílek et al. (2020) reports 15% of tidal features, including 5$\pm$2for streams and 5$\pm$2shells (which we consider together in our curve) at a SB limit of 28.5 - 29 mag $arcsec^{-}2$ for the $g$-band. Note that the $g$-band in the DES sample is fainter in average as the $r$-band by $\sim$ 0.5 mag $arcsec^{-}2$.
  • Sola et al. (2022) have characterised the morphology of more than 350 low surface brightness structures up to a distance of 42 Mpc through annotation of images from the Canada–France Imaging Survey (CFIS2) and the Mass Assembly of early-Type gaLAxies with their fine Structures survey (MATLAS3). They obtained 84 annotations for streams and 260 for shells, but out of these figures alone it is not possible to derive detectability figures for the brightest streams and shells, as there could be several of them in one image.
  • Miró-Carretero et al. (2023) reports a frequency of 12.2 $\pm$ 2.4% for Milky Way-like galaxies at a SB limit of 28.40 mag $arcsec^{-2}$.
  • Rutherford et al. (2024) use deep imaging from the Subaru-Hyper Suprime Cam Wide data to search for tidal features in massive [$log_{10}M_{\star}/M_{\odot}>10$] early-type galaxies (ETGs) in the SAMI Galaxy Survey. They report a tidal feature detection rate of 31 $\pm$2 % at a surface brightness limit of 27 $\pm$0.5 mag $arcsec^{-2}$ for the $r$-band. They calculate the surface brightness limit following the same approach as Atkinson et al. (2013), thus different to our method [$3\sigma$, 100 $arcsec^{2}$]. For comparison, our method applied to a test Subaru HSC image yields a surface brightness limit of 29.79 mag $arcsec^{-2}$.
  • In Skryabina et al. (2024) the results of visual inspection of a sample of 838 edge-on galaxies using images from three surveys: SDSS Strip-82, Subaru HSC and DESI (DECals, MzLS,BASS) are presented. In total, 49 tidal features out of 838 images are reported, equivalent to a frequency of $5.8\%$ at a SB limit of 28.60 mag $arcsec^{-}2$. In that study, tidal features include also disc deformations and tidal tails, typical of major mergers.
@@
  • Martin et al. (2022) reports on detection of tidal features inspecting mock-images produced using the NEWHORIZON cosmological simulations. Through production of surface brightness maps at different surface brightness limits, they predict the fraction of tidal features that can be expected to be detected at different limiting surface brightnesses. In this study, tidal features comprise: (i) Stellar streams, (ii) Tidal tails, (iii) Asymmetric stellar halos, (iv) Shells (v) Tidal bridges, (vi) Merger remnants (vii) Double nuclei, of which only (i) and (iv) are clearly of accreted nature. For a surface brightness limit of 35 mag arcsec-2), expert classifiers were able to identify specific tidal features in close to 100 per cent of galaxies ($M_{\star}$ $10^{9.5}M_{\odot}$), in agreement with our results. For the range of stellar mass of the DES sample, $log_{10}M_{\star}=10$, they report $\sim$ 25% detection rate for streams and shells together at a surface brightness limit of 30 mag $arcsec^{-}2$ in the $r$-band. The detection rate becomes 35% for a surface brightness limit of 31 mag $arcsec^{-}2$. These figures match very well our predictions with Auriga and IllustrisTNG50.
@@
  • In Khalid et al. (2024), the results of identification and classification of tidal featurest in LSST-like mock-images from cosmological simulations are reported. Four sets of hydrodynamical cosmological simulations are used (NewHorizons, EAGLE, IllustrisTNG and Magneticum). The frequency of tidal features, expressed in fractions of the total number of images, is between 0.32 and 0.40 showing consistency across the different simulations. Tidal Features comprise streams/tails, shells, Plumes or Asymmetric Stellar Halos and Double Nuclei. Looking only at streams and shells, the percentage of detections varies between 5-15% depending on the level of confidence, for a SB limit of 30.3 mag $arcsec^{-}2$ in the $r$-band.
@@
  • In Vera-Casanova et al. (2022), the authors report the results of inspecting surface brightness maps generated from 30 Auriga cosmological simulations (Grand et al., 2017) of Milky Way-like galaxies looking for the brightest streams. These simulations are the same we have applied in this work. They report that no streams have been detected in images with a surface brightness limit brighter than 25 mag arcsec-2 and that the stream detection frequency increases significantly between 28 and 29 mag arcsec-2. For a surface brightness limit of 28.50 mag $arcsec^{-}2$ in the $r$-band they achieve a detection rate of $\sim$ 18 - 30%. This is far higher than our results show with the same Auriga simulations and we believe the reason is the different pixel size of the mock-images. While we use the pixel size of the DECam instrument, 0.262 arcsec, equivalent to 0.08 kpc at 70 Mpc distance, in Vera-Casanova et al. (2022) they use a pixel size of 1 kpc. This is relevant for the visual inspection of count images, as it has an impact on the S/N per pixel and could account for a significant difference in the surface brightness limit at which the streams are detected.

### v2

@@ 5 Stream Detection
Figure 12: Detection rate curves for the COCO, TNG50 and Auriga cosmological simulations and the detection rate observed for the DES sample (black square indicates the dispersion in surface brightness limit of the DES sample and the dispersion in the detection rate according to a binomial distribution, see Miro-Carretero et al. (2024)). The percentage of the streams detected by visual inspection is plotted versus the surface brightness limit at which such detection was possible. The bars indicate the Poisson confidence interval.
@@ 5.1 Stellar Streams in the COCO Simulation
We have generated mock images with surface brightness limit between 25 and 34 mag arcsec-2 in intervals of 1 mag arcsec-2. As the depth of the image increases, that is, the fainter the surface brightness limit is, the clearer the underlying streams appear to the observer carrying out the visual inspection and facilitating their detection, as can be seen in Figure 11.
@@
The result is a curve indicating the percentage of streams detected within the COCO halo sample as a function of the image depth, that is, as a function of the image surface brightness limit. This is depicted in Figure 12. The curve follows an approximately linear trend between surface brightness limits of 26 and 34 mag arcsec-2, with a $12.5\%$ increase in streams detected per 1 mag arcsec-2 increase in surface brightness limit. In $97\%$ of the mock-images, we detect by visual inspection what we consider to be (part of) a stream up to a SB-limit of 34 mag arcsec-2. However, we estimate that only 89% of such faint structures could actually be measured with a reasonable level of error in real images.
Regarding the stream morphology, in the COCO sample, at SB-limit 28-29 arcsec-2, around 70-80% of the detected streams are shells –a segment of the wider morphology class known as umbrella, see stream morphology classification in Miro-Carretero et al. (2024)– while at SB-limit 34 arcsec-2 around 80-90% of the detected streams are shells, with only 2-3% of the streams displaying a circular morphology.
@@
Figure 14: Examples of TNG50 mock-images TNG50-394621, TNG50-440407, TNG50-469487 and TNG50-472548 at different surface brightness limits. From left to right 29, 30 and 31 mag arcsec-2. The streams appear clearer as the depth of the image increases from left to right. All images are $10\times 10$ arcmin.
5.2 Stellar Streams in the TNG50 Simulation
@@
We have transformed the surface brightness maps into images with counts and added sky background corresponding to 10 levels of surface brightness limit between 25 and 34 mag arcsec-2 (see Section 4). As we increase the depth of the image (apply a lower background noise level to the image, corresponding to a fainter surface brightness limit), streams become more visible and can be detected by visual inspection, as can be seen in Figure 14. We have visually inspected the resulting $\sim 600$ images and for each halo / host galaxy, the surface brightness limit at which a streams could first be visually appreciated was identified and noted. As with our analysis of COCO, FITS images were inspected with DS9.
Figure 15: Stellar streams detection rate as a function of the image surface brightness limit for the TNG50 simulation. The percentage of the streams detected by visual inspection is plotted versus the surface brightness limit at which such detection was possible. The solid line corresponds to a stellar mass range of $\log_{10}M_{\star}/\mathrm{M_{\odot}}$ = 10.48 and $\log_{10}M_{\star}/\mathrm{M_{\odot}}$ = 10.76, while the dashed line corresponds to an extended mass range up to $\log_{10}M_{\star}/\mathrm{M_{\odot}}$ = 10.9.
@@ 5.3 Stellar Streams in the Auriga Simulation
We have available 30 Auriga zoom simulations of MW-mass halos at $z<0.02$. Since this sample may not be statistically representative, and in order to compare with the COCO and TNG50 simulations, three surface brightness maps have been generated from each halo, taking the three orthogonal axes of the simulation coordinate system as lines of sight. None of the axes are aligned with face-on or edge-on directions so that overall the orientation of the galaxies is random. This process results in 90 images, each of which we combine with a range of sky backgrounds between the SB limit of 25 and 34 mag arcsec-2, at intervals of 1 mag arcsec-2, thus yielding 900 mock images in total. Examples of halos with streams of different morphology can be seen in Figure 16.
@@
Figure 18: Comparison of the stellar stream detection rate as a function of the image surface brightness limit for TNG50 (red), COCO (blue) and AURIGA (green) with the results of observations. (1) The calculation of the SB limit follows a different method (2) The detection rate has been derived for all LSB features, including streams.
Figure 12 shows the detection rate curve obtained for Auriga from visual inspection of the mock images, together with those obtained for COCO and TNG50, and the observational results of the DES sample. The figure shows the percentage of the sample for which at least one stream is detected for each of the 10 surface brightness limit levels analysed. The curve corresponds to detections for which there is a reasonable level of confidence that are streams. When we detect a LSB feature but have doubts about whether that LSB feature constitutes a stream, we take the conservative approach of not counting it towards the results included in the figure. Three of the halos (Au11 and Au30) seem to show an ongoing major merger between two galaxies of similar apparent size, clearly visible at a surface brightness limit of 26 and 27 mag arcsec-2, respectively. These three halos have been discarded, because we are looking for remnants of minor mergers only. LSB structures that could appear similar to streams at fainter surface brightness limits are not accounted for in the detectability curve.
@@
Table 2: Percentages of the principal stream morphologies in observation and simulations.
@@
Note. Percentages the principal stream morphologies shell and circle identified in the the DES sample and in the mock images for the COCO, TNG50 and AURIGA simulations.The 4th column indicates the total number of streams detected in the sample. Note that some images present more than one stream.
6 Stream Photometry
We have carried out a photometric analysis of the mock images generated from cosmological simulations though the process described in Section 4. We present here the results of the characterisation and photometric analysis of the mock-images generated from the COCO and TNG50 simulations and compare them with the ones obtained from the DES observations sample reported in (Miro-Carretero et al., 2024)). In this paper, we restrict this photometric analysis to the two cosmological volume simulations, leaving a similar study of the more detailed higher-resolution Auriga simulations for future work (note also that due to the smaller number of Auriga simulations, such an analysis would have less statistical significance). To measure photometry parameters such as surface brightness and colours on the mock images the corresponding surface brightness maps have been generated for the SDSS $r$,$g$ and $z$ filters and transformed into counts images.
In order to compare the photometry of images with a similar surface brightness limit, the output of the simulations has been enhanced by superimposing a real DES sky background selected to have a surface brightness limit corresponding to the average surface brightness limit of the DES image sample with streams, namely 28.65 mag arcsec-2 (see Section 4. As a result, only those mock images have been selected for the comparison, in which a stream was detected at a surface brightness limit of $\leq$ 28.65 This results in 21 mock-images from the COCO simulation and 8 mock images from the TNG50 simulation. Their photometry parameters are compared against those from 63 images of the DES sample with streams. This means that a statistically sound comparison is not possible, nevertheless the comparison is useful to see whether the cosmological simulations can reproduce real photometric observations, which requires realistic predictions of particle metallicity and age, which in turn is the result of the modelling of the physical processes involved.
The comparison between the mock-images photometry and the observations from the DES galaxy sample is carried out on the basis of the average stream surface brightness, average stream $(g-r)_{0}$ colour as well as the average distance of stream to the host centre. The resulting histograms are shown in Figures 19, 20 and 21.
@@ 7 Discussion
We now compare the predictions of the cosmological simulations with one another and with the observational data regarding frequency, detectability and morphology of streams as well as their photometric properties. Looking at Figure 18, the cosmological simulations all seem to predict that, for a surface brightness limit of 32 $\mathrm{mag\,arcsec^{-2}}$, almost 70% of galaxies in the mass range we study have one or more detectable streams.
However, the simulations show discrepancies with one another regarding detection rates for surface brightness limits brighter than $32\,\mathrm{mag\,arcsec^{-2}}$. In particular, the two hydrodynamical simulations, TNG50 and Auriga, agree well with each other but predict a lower detectability rate than COCO. For limits fainter than $32\,\mathrm{mag\,arcsec^{-2}}$, however, TNG50 has a lower detection rate than COCO and Auriga (which agree well with each other, in this regime). The simulations also show discrepancies with one another regarding the morphology of the streams detected, as will be discussed further down in this section.
@@
As discussed in Cooper et al. (2017), when comparing particle tagging simulations with hydrodynamical methods, even with identical initial conditions, it can be difficult to separate effects due to the dynamical approximation of particle tagging from the effects of different star formation models. In our case, the IllustrisTNG and COCO models have both been calibrated to observations of the galaxy mass and luminosity functions at $z=0$ (in the case of COCO via the semi-analytic model of Lacey et al., 2016), as well as other low-redshift data, notably the galaxy size–mass relation. The fundamental relationship between stellar mass and virial mass in both simulations agrees well with (for example) that inferred from galaxy abundance matching. Although the typical star formation histories of host galaxies in our sample, and their stream progenitors, may still differ in detail between the two simulations, we expect that they are broadly similar. This makes it more likely (although by no means certain) that the differences we observe are related to dynamical factors, rather than differences in how stars populate dark matter halos.
@@
Cooper et al. (2017) explore the consequences of neglecting these factors when predicting satellite disruption (and hence stream and stellar halo formation) in particle tagging models. In the hydrodynamical model used in that work, massive satellites were found to disrupt somewhat earlier than realizations of the same systems in a particle tagging model [^7] This of course depends on the detail of the hydrodynamical scheme and its subgrid recipes for star formation and feedback; the specific scheme used in IllustrisTNG was not examined by Cooper et al. (2017). A detailed case study of a single bright stream relevant to this discussion is given in Appendix A of that paper.. This could explain why more streams are visible at surface brightness limits of $\lesssim 28$ mag arcsec-2 in COCO compared to IllustrisTNG; if satellites are disrupted earlier in IllustrisTNG, their streams may have more time to phase mix, lowering their surface brightness. A similar argument could also explain the greater abundance of shells detected in COCO. Shells originate from satellites on radial orbits (Newberg & Carlin, 2016). The relative fraction of radial and circular orbits may differ between IllustrisTNG and COCO, because the evolution of the progenitor’s orbit during pericentric passages (due to exchange of angular momentum with the host) may be significantly different with and without a massive baryonic disk. A related effect is explored by Valenzuela & Remus (2024), who study the relation between the rotation of the hosts and the presence and morphology of tidal features One of the conclusions of that work is that shells appear more frequently in slow rotating hosts than in fast rotating hosts. This hypothesis could be tested in future work. More generally, further exploration of these differences particle tagging and hydrodynamical simulations could help to understand how observations of streams can constrain the nature of the gravitational potentials of stream progenitors and their host galaxies.
@@
Turning now to comparison between the simulations and the observational data, $8.7\pm 1.1\%$ of galaxies in the DES sample have detectable streams at an average $r$-band surface brightness limit of 28.65 mag arcsec-2. This seems to match very well with the predictions of TNG50 and Auriga, and is lower than the prediction from COCO.
@@
Stream morphology is not an observable that can be used reliably to constrain simulations, because it is strongly dependent on the line of sight along which the stream is observed. Nevertheless, comparison of the fraction of streams with different morphologies is interesting because it could be related to the same dynamical differences between simulations methods that give rise to differences in stream abundance. 20-43% of streams in the TNG50 sample are shells, which is within the range found in the DES sample, and 10-16% have circular morphology, which is also not too far from the observations. As noted above, there is a significant discrepancy in the relative fractions of different stream morphologies between observations and the COCO simulations. In the COCO mock images, 70-90% of streams are identified as having Umbrella/Shell morphology, while in the DES sample, only 27-38% of streams are identified as such. The COCO mock images show only 2-3 % of streams with circular morphologies, compared to 21-35 % in the DES images. Auriga predicts 41-62% of streams with Umbrella/Shell morphologies, significantly above the fraction in the DES sample, and 10-24% of stream with circular morphology, somewhat below the DES figures. Table 2 summarises the stream morphology findings for the different samples derived from the total number of streams detected. Note that more than one stream is detected in some of the halos. It is beyond the scope of this work to analyse the reasons for these discrepancies in stream morphology; as noted above, further work would be a valuable contribution to understanding the relationship between the observable properties of streams and the galaxy formation process as a whole.
Regarding the stream photometry measurements, the stream average surface brightness range in the COCO simulation matches generally well the observations in the DES stream sample, though the distribution is skewed towards the brighter end of the range (Figure 19). This is consistent with the abundance of shell-shaped streams, typically brighter than other stream morphologies, as confirmed by the observations. The range of the stream average surface brightness in the TNG50 simulation falls pretty much within the central region of the DES sample range.
The stream $(g-r)_{0}$ colour distribution of the COCO simulations also matches generally well the observations. The mean value of the COCO simulation distribution is 0.54$\pm$0.12 mag , versus 0.57$\pm$0.14 mag for the DES observations sample. However, the simulations are slightly skewed towards the blue end as can be seen in Figure 20. The TNG50 simulation, though not statistically significant, as we have only the value for 8 streams, shows a $(g-r)_{0}$ colour distribution spanning almost the full range covered by the DES observations.
Consistent with the overabundance of shell stream morphology in the COCO simulations, the average distance of the streams to the host galaxy is between 10 and 30 kpc, thus covering the lower end of the DES sample distance range, see Figure 21. The streams from the TNG50 simulation are at an average distance between 20 and 50 kpc from the host centre, covering the central part of the distance range observed in the DES sample.
@@ 7.1 Comparison with Previous Work
  • Miskolczi et al. (2011) reports a frequency of 6-19% at a SB-limit of $\sim$ 27 mag arcsec-2. We consider the lowest value of 6% as the most reliable for maximising purity in the detection of streams (see Figure 18).
@@
  • Morales et al. (2018) reports a frequency of $\sim$ 10% at a SB limit of 28 mag arcsec-2.
  • Bílek et al. (2020) reports 15% of tidal features, including 5$\pm$2for streams and 5$\pm$2shells (which we consider together in our curve) at a SB limit of 28.5 - 29 mag arcsec-2 for the $g$-band. Note that the $g$-band in the DES sample is fainter in average as the $r$-band by $\sim$ 0.5 mag arcsec-2.
  • Sola et al. (2022) have characterised the morphology of more than 350 low surface brightness structures up to a distance of 42 Mpc through annotation of images from the Canada–France Imaging Survey (CFIS2) and the Mass Assembly of early-Type gaLAxies with their fine Structures survey (MATLAS, Duc, 2020; Bílek et al., 2020). They obtained 84 annotations for streams and 260 for shells, but out of these figures alone it is not possible to derive detectability figures for the brightest streams and shells, as there could be several of them in one image.
  • Miró-Carretero et al. (2023) report on the search of stellar tidal streams in DESI Legacy Survey images of MW-like galaxies, at distances between 25 and 40 Mpc, from the SAGA II galaxy sample (Geha et al., 2017; Mao et al., 2021). Applying the same detection and characterisation methods as in this work, their statistical analysis yields a stream detection frequency of 12.2 $\pm$ 2.4% at a SB limit of 28.40 mag arcsec-2.
  • Rutherford et al. (2024) use deep imaging from the Subaru-Hyper Suprime Cam Wide data to search for tidal features in massive [$\log_{10}M_{\star}/M_{\odot}>10$] early-type galaxies (ETGs) in the SAMI Galaxy Survey. They report a tidal feature detection rate of 31 $\pm$2 % at a surface brightness limit of 27 $\pm$0.5 mag arcsec-2 for the $r$-band. They calculate the surface brightness limit following the same approach as Atkinson et al. (2013), thus different to our method [$3\sigma$, 100 $arcsec^{2}$]. For comparison, our method applied to a test Subaru HSC image yields a surface brightness limit of 29.79 mag arcsec-2.
  • In Skryabina et al. (2024) the results of visual inspection of a sample of 838 edge-on galaxies using images from three surveys: SDSS Strip-82, Subaru HSC and DESI (DECals, MzLS,BASS) are presented. In total, 49 tidal features out of 838 images are reported, equivalent to a frequency of $5.8\%$ at a SB limit of 28.60 mag arcsec-2. In that study, tidal features include also disc deformations and tidal tails, typical of major mergers.
@@
  • Martin et al. (2022) reports on detection of tidal features inspecting mock-images produced using the NEWHORIZON cosmological simulations. Through production of surface brightness maps at different surface brightness limits, they predict the fraction of tidal features that can be expected to be detected at different limiting surface brightnesses. In this study, tidal features comprise: (i) Stellar streams, (ii) Tidal tails, (iii) Asymmetric stellar halos, (iv) Shells (v) Tidal bridges, (vi) Merger remnants (vii) Double nuclei, of which only (i) and (iv) are clearly of accreted nature. For a surface brightness limit of 35 mag arcsec-2), expert classifiers were able to identify specific tidal features in close to 100 per cent of galaxies ($M_{\star}$ $10^{9.5}M_{\odot}$), in agreement with our results. For the range of stellar mass of the DES sample, $\log_{10}M_{\star}/M_{\odot}=10$, they report $\sim$ 25% detection rate for streams and shells together at a surface brightness limit of 30 mag arcsec-2 in the $r$-band. The detection rate becomes 35% for a surface brightness limit of 31 mag arcsec-2. These figures match very well our predictions with Auriga and TNG50.
@@
  • In Khalid et al. (2024), the results of identification and classification of tidal featurest in LSST-like mock-images from cosmological simulations are reported. Four sets of hydrodynamical cosmological simulations are used (NewHorizons, EAGLE, IllustrisTNG and Magneticum). The frequency of tidal features, expressed in fractions of the total number of images, is between 0.32 and 0.40 showing consistency across the different simulations. Tidal Features comprise streams/tails, shells, Plumes or Asymmetric Stellar Halos and Double Nuclei. Looking only at streams and shells, the percentage of detections varies between 5-15% depending on the level of confidence, for a SB limit of 30.3 mag arcsec-2 in the $r$-band.
@@
  • In Vera-Casanova et al. (2022), the authors report the results of inspecting surface brightness maps generated from 30 Auriga cosmological simulations (Grand et al., 2017) of MW-like galaxies looking for the brightest streams. These simulations are the same we have applied in this work. They report that no streams have been detected in images with a surface brightness limit brighter than 25 mag arcsec-2 and that the stream detection frequency increases significantly between 28 and 29 mag arcsec-2. For a surface brightness limit of 28.50 mag arcsec-2 in the $r$-band they report a detection rate of $\sim$ 18 - 30%. This is far higher than our results show with the same Auriga simulations and we believe the reason is the different pixel size of the mock-images. While we use the pixel size of the DECam instrument, 0.262 arcsec, equivalent to 0.08 kpc at 70 Mpc distance, in Vera-Casanova et al. (2022) they use a pixel size of 1 kpc. This is relevant for the visual inspection of counts images, as it has an impact on the S/N per pixel and could account for a significant difference in the surface brightness limit at which the streams are detected. Rebinning the DECam images to a 1 kpc pixel size shows that this difference exceeds 2 mag arcsec-2.